OpenStudy (anonymous):

11/X = 20 Solve for X

OpenStudy (calculusxy):

\[\frac{ 11 }{ x }=\frac{ 20 }{ 1 }\] Now just cross-multiply and find the value for x.

OpenStudy (calculusxy):

\[11 \times 1 = 20 \times x\] \[11=20x\]

OpenStudy (anonymous):

thnx

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (anonymous):

.55 right?

OpenStudy (calculusxy):

Yeah